OpenStudy (anonymous):

GRAPH THE LINE CONTAINING THE GIVEN PAIR OF POINTS AND FIND THE SLOPE (1,0) AND (6,-7)

OpenStudy (anonymous):

-7/6 is the incorrect answer

OpenStudy (saifoo.khan):

Oh,..thats the slope.

OpenStudy (saifoo.khan):

y = mx + c 0 = -7/6 + c c = 7/6

OpenStudy (saifoo.khan):

y = mx + c 6y = -7x + 7

OpenStudy (saifoo.khan):

this is your eq.

OpenStudy (dumbcow):

m=slope = rise/run = (y2-y1)/(x2-x1) = (-7-0)/(6-1) = -7/5 point slope form: y-y1 = m(x-x1) y-0 = (-7/5)(x -1) y = -7/5x + 7/5

OpenStudy (saifoo.khan):

Sorry my bad. it was 6-1 = 5 :/

OpenStudy (saifoo.khan):

rest all is correct. :D